A cylinder can has a diameter of 20cm and a height of 30cm. How many litres does it hold? (1ℓ = 10³ cm³)
Answer by maheswari2188(4)

Volume of a cylinder = πr^2h, where r is the diameter, h is the height of the cylinder.
V = π * 20 * 20 * 30
= 3.14159265 * 20 * 20 * 30
= 37699.1118 cm^3
= 3.77 Litres